# Area Selective Chemical Vapor Deposition of Gold by Electron Beam Seeding Zusammenfassung

### v1

@@ Acknowledgements
Gold particles are observed only for the highest dose as evidenced by the close-up high resolution SE micrograph provided in Figure SI.2(c), which is a magnified view of the highlighted region in Figure SI.2(b). This implies that under high-vacuum (HV) conditions, surface activation by electrons, as also demonstrated by Walz et al. 13, may be achievable. Dark contrast in squares may arise from hydrocarbon contamination on the surface caused by residual gas composition in high vacuum 52. These deposits are commonly referred to as carbonaceous contamination.
Thickness evaluation
@@
Atomic force microscopy (AFM) measurements were carried out to determine the thickness of the autocatalytically grown gold films. An SE image of the autocatalytically grown films at an elevated temperature of 80 °C for 30 min is shown in Figure SI.5(a). Seeding for the films was done at 80 °C. Figure SI.5(b) shows the AFM thickness profiles of the squares in the 3rd and 6th rows, which were seeded with a minimum pitch of 5 nm. The thickness of the film versus the deposition time of the seed layer for each individual square is displayed in Figure SI.5(c). The data was compared to an exponential asymptotic function that exhibits a saturation of thickness. The saturation in autocatalytic CVD gold growth was observed on certain seed squares with a deposition time of 5.6 s for a size of 2 µm x 2 µm, corresponding to an electron dose of 140 pC/µm or about 875 electrons per nm. Corresponding thickness of the film autocatalytically grown for 30 min on this seed layer is about 6.5 nm. Gold films after seeding and autocatalytic growth at 80 °C for longer times of 60 and 90 mins were additionally investigated. A saturation in achievable film thickness in both cases was also observed as discussed in the main manuscript. Autocatalytic growth for 60 min only resulted in additional 0.7 nm, and even 90 minutes total growth time resulted only in 8.5 nm.
Transmission Electron Microscopy
@@
For transmission electron microscopy (TEM) investigations, two different samples were used. Gold films were grown autocatalytically on substrates at a temperature of 90 °C for 30 min, but were seeded at 30 °C for sample (a) and 90 °C for sample (b) (Figure SI.6(a) and (b) respectively). Gold films were grown on a n-type (100) Si chip with native oxide for sample (a) and on a p-type (100) Si chip with native oxide for sample (b). TEM lamellae were prepared according to the standard lift-out protocol using dual beam FIB-SEM Tescan Lyra3. The TEM images were captured using a probe-corrected ThermoFisher Scientific Titan Themis 200 G3. Figure SI.6 shows the full STEM micrographs used for the main manuscript. In both instances, periodic atomic order of specific gold grains can be observed.
Monte-Carlo simulations
@@
The energy deposited by electrons in the material was extracted using a Monte-Carlo method proposed by David C. Joy’s single scattering model 53. The analytical Mott expression 54 implemented the inelastic scattering event, and the modified Bethe expression 53 was used for the elastic energy loss, which is an average energy loss by an electron due to all inelastic scattering events. The structure representing the FEBID-grown pillar was implemented with a cylinder and a truncated cone above the cylinder. Trajectory and energy loss of each electron were recorded and used to calculate the total energy lost by primary electrons on average. Figure SI.7 depicts the paths of 500 electrons with 15 keV energy, traveling within an object that has a top truncated cone with a 10 nm radius and a height of 150 nm, as well as a cylinder with a 25 nm radius. The object has a composition ratio of Au:C as 0.28:0.72, with a density of 6.5 g/cm. The energy transfer value for 4 g/cm lower bound was calculated using the same method. The object’s geometry was derived from SEM images. The energy deposited per each primary electron and line segment between two collisions at the apex of the pillar is shown in the inset of the figure.
Temperature Distribution in Pillars Induced by the Electron Beam
@@
Convection losses in high vacuum ( mbar) are negligible because the mean free path of the residual gas molecules is greater than the size of the vacuum chamber. The radiative losses can be estimated by applying the Stefan-Boltzmann law, where the total radiative power is the difference between the power emitted by the column and the power absorbed by the surrounding medium: , where is the surface area of the column, W/mK is the Stefan-Boltzmann constant, is the emissivity, and is the ambient temperature of the vacuum chamber. The typical pillar deposited at 30 °C has a surface area of m with a radius of nm and a length of nm. In order to estimate the maximum radiative losses, we assume a maximum emissivity of the deposit of 1. For the aforementioned reason, we ignore any irregularities in temperature distribution and assume uniform heating of the entire pillar, with its length being equal to the previously determined temperature of 112 °C (385 K). In this instance, the maximum power output is W. As per Monte-Carlo simulations, a primary electron of 15 kV conveys at least 0.75 kV of its initial energy to the pillar. This then results in an implanted power of W by a focused electron beam with a current of 100 pA. If the power transfer from the focused electron beam is compared with the maximum estimated radiative loss, the latter is at least 2 orders of magnitude lower and thus negligible, which justifies a linear temperature gradient in the pillar.
@@
The following schematics are proposed for a better understanding of the heat transfer and temperature distribution in the electron beam grown pillars. The left-hand side of Figure SI.8(a) depicts the time evolution of a single pillar.
@@
Let us first consider the stage temperature T set at the room temperature of 30 °C. At smaller sizes, the majority of the heat supplied via electron beam is dissipated within the substrate. However, once the pillar reaches a certain height that corresponds to the stopping range of primary electrons in the rod, the implanted energy from primary electrons into the pillar will be maximised. The temperature of the top of the pillar exceeding the base temperature (i.e. the temperature of the substrate) generates an almost linear temperature gradient along the axis of the pillar. This can be attributed to a quasi-continuous heat supply by the electron beam and the stage temperature, T, remaining constant, resulting in a steady solution of the heat transfer equation (cf. 55). This occurs because there are no significant convectional or radiative losses as estimated previously. As a result, the temperature of the upper portion of the pillar may reach the threshold for the autocatalytic growth of gold particles, which only becomes visible once the gold crystals increase to a size that can be observed by SEM imaging at the sidewall of the pillar. During the experiment with Au(acac)Me, this occurred at about 2 min 30 s with a 15 keV 100 pA electron beam. The temperature distribution within the upper section of the pillar remains uncertain because of the presence of gold crystals, which impact heat conductivity.
@@
If a pillar is now deposited with the same deposition parameters and total time, but on the stage at an elevated temperature of 70 °C, an autocatalytic process will take place at a lower point along the pillar axis (see Figure SI.8(a)-(iv)). By progressively increasing the stage temperature, gold crystals will eventually appear on the pillar alongside the sample surface, as shown in Figure SI.8(a)-(v). This enables the calculation of T - the temperature induced by the electron beam at the point along the pillar where the CVD process first occurs (see Figure SI.8(a)). T remains constant during the specified deposition period. If we plot all temperature distributions for each pillar grown at different T (refer to Figure SI.8(b)), we can observe that the point on the pillar heated to the temperature T shifts linearly to the sample surface and collapses with the surface at around 112 °C.

### v2

@@ 4 Control Experiments
The first control test of pattern array deposition was performed without the Au(acac)Me2 precursor gas being introduced into the vacuum chamber. The array pattern was irradiated with the electron beam while residual chamber gas was present at a base pressure of mbar, both at room temperature (30 °C) and at 90 °C. Subsequently, both irradiated areas were left at elevated temperature of 90 °C for 50 min with the gold precursor gas flow enabled. SE micrographs presenting the results are displayed in Figure SI.9(a) and (b). Gold particles are observed only for the highest dose as evidenced by the close-up high resolution SE micrograph provided in Figure SI.6(c), which is a magnified view of the highlighted region in Figure SI.6(b). This implies that under high-vacuum (HV) conditions, surface activation by electrons, as also demonstrated by Walz et al. [13], may be achievable. Dark contrast in squares may arise from hydrocarbon contamination on the surface caused by residual gas composition in high vacuum [54]. These deposits are commonly referred to as carbonaceous contamination.
The second control experiment was done with the seeding by carbon (C10H8 - naphthalene) and tungsten (W(CO)6 - tungsten hexacarbonyl) precursors at 90 °C substrate temperature. Deposited arrays were heated to 90 °C, while being exposed to the gold precursor for 30 min. The results are shown in Fig. SI.10, and depict no significant amount of gold grown after the experiment in both cases.
The last control experiment was done with seeding by Au(acac)Me2 precursor at 30 and 80 °C substrate temperature. Test heating without precursor gas flow was then done at 80 °C for 30 min and then directly compared to as-deposited seed layers at corresponding temperatures. The results are shown in Fig. SI.11, and depict no gold growth on the seeds at both temperatures in comparison to the as-deposited seed layers.
5 Thickness Evaluation
@@
Atomic force microscopy (AFM) measurements were carried out to determine the thickness of the autocatalytically grown gold films. An SE image of the autocatalytically grown films at an elevated temperature of 80 °C for 30 min is shown in Figure SI.12(a). Seeding for the films was done at 80 °C. Figure SI.12(b) shows the AFM thickness profiles of the squares in the 3rd and 6th rows, which were seeded with a minimum pitch of 5 nm. The thickness of the film versus the deposition time of the seed layer for each individual square is displayed in Figure SI.12(c). The data was compared to an exponential asymptotic function that exhibits a saturation of thickness. The saturation in autocatalytic CVD gold growth was observed on certain seed squares with a deposition time of 5.6 s for a size of 2 µm x 2 µm, corresponding to an electron dose of 140 pC/µm2 or about 875 electrons per nm2. Corresponding thickness of such a film is about 6.5 nm. Gold films after seeding and autocatalytic growth at 80 °C for longer times of 60 and 90 mins were additionally investigated. A saturation in achievable film thickness in both cases was also observed as discussed in the main manuscript. Autocatalytic growth for 60 min only resulted in additional 0.7 nm, and even 90 minutes total growth time resulted only in 8.5 nm.
Additionally, a gold seed layer was investigated in AFM, which was deposited at 90 °C substrate temperature. SE micrographs as well as AFM profiles are depicted in Fig. SI.13.
6 Transmission Electron Microscopy
@@
For transmission electron microscopy (TEM) investigations, two different samples were used. Gold films were grown autocatalytically on substrates at a temperature of 90 °C for 30 min, but were seeded at 30 °C for sample (a) and 90 °C for sample (b) (Figure SI.14(a) and (b) respectively). Gold films were grown on a n-type (100) Si chip with native oxide for sample (a) and on a p-type (100) Si chip with native oxide for sample (b). TEM lamellae were prepared according to the standard lift-out protocol using dual beam FIB-SEM Tescan Lyra3. The TEM images were captured using a probe-corrected ThermoFisher Scientific Titan Themis 200 G3. Figure SI.14 shows the full STEM micrographs used for the main manuscript. In both instances, periodic atomic order of specific gold grains can be observed.
Additionally, nanopillars were grown on a standard FIB-TEM Cu grid to investigate crystallinity of gold grains, described in Section 3 of the main manuscript. Fig. SI.14(c) contains a HAADF TEM image of one pillar together with a close-up images from highlighted area (d) with gold crystals (grown autocatalytically) and (e) FEBID grown pillar. Due to the mechanical vibrations, it is impossible to resolve periodic atomic pattern within crystals.
7 Monte-Carlo Simulations
@@
The energy deposited by electrons in the material was extracted using a Monte-Carlo method proposed by David C. Joy’s single scattering model [55]. The analytical Mott expression [56] implemented the inelastic scattering event, and the modified Bethe expression [55] was used for the elastic energy loss, which is an average energy loss by an electron due to all inelastic scattering events. The structure representing the FEBID-grown pillar was implemented with a cylinder and a truncated cone above the cylinder. Trajectory and energy loss of each electron were recorded and used to calculate the total energy lost by primary electrons on average. Figure SI.15 depicts the paths of 500 electrons with 15 keV energy, traveling within an object that has a top truncated cone with a 10 nm radius and a height of 150 nm, as well as a cylinder with a 25 nm radius. The object has a composition ratio of Au:C as 0.28:0.72, with a density of 6.5 g/cm3. The energy transfer value for 4 g/cm3 lower bound was calculated using the same method. The object’s geometry was derived from SEM images. The energy deposited per each primary electron and line segment between two collisions at the apex of the pillar is shown in the inset of the figure.
8 Temperature Distribution in Pillars Induced by the Electron Beam
@@
Convection losses in high vacuum ( mbar) are negligible because the mean free path of the residual gas molecules is greater than the size of the vacuum chamber. The radiative losses can be estimated by applying the Stefan-Boltzmann law, where the total radiative power is the difference between the power emitted by the column and the power absorbed by the surrounding medium: , where is the surface area of the column, W/m2K4 is the Stefan-Boltzmann constant, is the emissivity, and is the ambient temperature of the vacuum chamber. The typical pillar deposited at 30 °C has a surface area of m2 with a radius of nm and a length of nm. In order to estimate the maximum radiative losses, we assume a maximum emissivity of the deposit of 1. For the aforementioned reason, we ignore any irregularities in temperature distribution and assume uniform heating of the entire pillar, with its length being equal to the previously determined temperature of 112 °C (385 K). In this instance, the maximum power output is W. As per Monte-Carlo simulations, a primary electron of 15 kV conveys at least 0.75 kV of its initial energy to the pillar. This then results in an implanted power of W by a focused electron beam with a current of 100 pA. If the power transfer from the focused electron beam is compared with the maximum estimated radiative loss, the latter is at least 2 orders of magnitude lower and thus negligible, which justifies a linear temperature gradient in the pillar.
@@
The following schematics are proposed for a better understanding of the heat transfer and temperature distribution in the electron beam grown pillars. The left-hand side of Figure SI.16(a) depicts the time evolution of a single pillar.
@@
Let us first consider the stage temperature Ts set at the room temperature of 30 °C. At smaller sizes, the majority of the heat supplied via electron beam is dissipated within the substrate. However, once the pillar reaches a certain height that corresponds to the stopping range of primary electrons in the rod, the implanted energy from primary electrons into the pillar will be maximised. The temperature of the top of the pillar exceeding the base temperature (i.e. the temperature of the substrate) generates an almost linear temperature gradient along the axis of the pillar. This can be attributed to a quasi-continuous heat supply by the electron beam and the stage temperature, Ts, remaining constant, resulting in a steady solution of the heat transfer equation (cf. [57]). This occurs because there are no significant convectional or radiative losses as estimated previously. As a result, the temperature of the upper portion of the pillar may reach the threshold for the autocatalytic growth of gold particles, which only becomes visible once the gold crystals increase to a size that can be observed by SEM imaging at the sidewall of the pillar. During the experiment with Au(acac)Me2, this occurred at about 2 min 30 s with a 15 keV 100 pA electron beam. The temperature distribution within the upper section of the pillar remains uncertain because of the presence of gold crystals, which impact heat conductivity.
@@
If a pillar is now deposited with the same deposition parameters and total time, but on the stage at an elevated temperature of 70 °C, an autocatalytic process will take place at a lower point along the pillar axis (see Figure SI.16(a)-(iv)). By progressively increasing the stage temperature, gold crystals will eventually appear on the pillar alongside the sample surface, as shown in Figure SI.16(a)-(v). This enables the calculation of Ti - the temperature induced by the electron beam at the point along the pillar where the CVD process first occurs (see Figure SI.16(a)). Ti remains constant during the specified deposition period. If we plot all temperature distributions for each pillar grown at different Ts (refer to Figure SI.16(b)), we can observe that the point on the pillar heated to the temperature Ti shifts linearly to the sample surface and collapses with the surface at around 112 °C.
9 Line width experiment
An additional line width experiment was conducted. The seed layer with a simple line pattern (as shown in Fig. SI.17 (a)) was deposited onto a 90 °C substrate, followed by a standard autocatalytic growth of gold at 90 °C. The results are presented in Fig. SI.17 (b). For 5 nm pitch patterning with 5 kV and 100 pA electron beam, a saturation in growth occurs at 80 repeats and reaches approximately 40 nm.
The actual electron beam probe size is much smaller (below 5 nm) than the obtained line width. The lateral structure widening is caused by SE and BSE and strongly depends on the acceleration voltage of the beam as known from e-beam lithography literature. The best area-selective fidelity could be achieved by setting the lowest available primary beam voltage to lower than 1 kV. Experimentally, at 5 kV we observed the broadening to be less than 100 nm, which can be seen in the inset (c) and (d) of Fig. 2.(b) of the main manuscript.